OpenStudy (anonymous):

FIND THE RATIO AND SIMPLIFY. FIND THE RATIO OF THE SHORTEST TO THE LONGEST 6, 11.8, AND 13.2

OpenStudy (kropot72):

Which are the shortest and the longest numbers?

OpenStudy (anonymous):

\[\frac{6}{13.2}=\frac{60}{132}=\frac{5}{11} \]